OpenStudy (anonymous):

The equation of a circle is (x + 4)2 + (y + 2)2 = 16. Determine the coordinates of the center of the circle and the length of the radius.

OpenStudy (anonymous):

(-4,-2) length of radius= 4 unit

Directrix (directrix):

(x + 4)^2 + (y + 2)^2 = 4^2 ((x + (-4) ) ^2 + (y + (+2))^2 = 4^2 is written in the general form of the equation of a circle. Center: (-4, -2) and radius: 4

Directrix (directrix):

Questions?

OpenStudy (anonymous):

thank you I understand!

Directrix (directrix):

Great. Just keep the general form of a circle in mind and you will be able to do these. :)